# Thermodynamic analysis of shift-symmetric black-hole spacetimes in Horndeski gravity

### v1

@@ IV General thermodynamical remarks
boundary of $\text{boundary of $X$}=$ $\bigg{(}-4\pi\beta\sqrt{h}N\delta X\Big{(}\sqrt{2}s\sqrt{-X}\left(f\left(24X^{2}F_{5X}+60F_{5}X-2XG_{5XX}-3G_{5X}\right)+r^{2}G_{3X}+G_{5X}\right)$
$+4\sqrt{f}r\left(2X\left(2XF_{4X}+G_{4XX}\right)+G_{4X}\right)+32\sqrt{f}F_{4}rX\Big{)}\bigg{)}\bigg{|}_{r_{h}}^{\infty}~{}.$ (48)
@@
boundary of $\text{boundary of $f$}=$ $\bigg{(}-4\pi\beta N\delta f\left(-2G_{4}r+2X\left(\sqrt{2}\sqrt{f}s\sqrt{-X}\left(12F_{5}X-G_{5X}\right)+4F_{4}rX+2rG_{4X}\right)\right)\frac{\sqrt{h}}{\sqrt{f}}\bigg{)}\bigg{|}_{r_{h}}^{\infty}~{}.$ (49)
@@
boundary of $\text{boundary of {\hbox{X}}}=\frac{4\pi\beta\delta X\sqrt{h}\left(fh^{\prime}-fh\right)\left(-2\sqrt{2}f^{3/2}s(-X)^{3/2}\left(12F_{5}X-G_{5X}\right)+G_{2}r^{3}+2G_{4}r\right)}{f^{3/2}X\left(rh^{\prime}+h\right)}~{}.$ (50)
Interestingly, for homogeneous metrics, which satisfy $h(r)=f(r)$, the first parenthesis vanishes, since $fh^{\prime}-fh=0$ so a boundary term for $X$ does not survive. In the non-homogeneous case, i.e. $h\neq f$, the boundary term of $X$ may vanish when the metric function $f$ satisfies
@@
This implies that a radially dependent scalar field cannot support primary-charged black hole solutions in the context of shift-symmetric Horndeski theories, due to the vanishing of the boundary term of $X$, when the spacetime is characterized by a homogeneous metric or when $f$ satisfies the aforementioned relation (51). In the case where the scalar field is linearly time-dependent, a nonzero conjugate momentum term arises when moving to the Hamiltonian formulation [107, 105], resulting to a non-zero boundary term for $X$. Hence, a linearly time-dependent scalar field is the only way that one can generate a black hole carrying primary scalar hair in the context of homogeneous metric, and the corresponding field has to have a non-zero conjugate momentum [31, 105], just like the case of electric fields in electromagnetism [112, 113, 85]. Regardless of the fact of the vanishing of the boundary term of $X$, a primary scalar hair has to be characterized by a non-trivial Noether current, which cannot be the case of purely radial scalar field, as it has been shown in [31].
@@ IV.1 Homogeneous Spacetimes
$\delta\mathcal{S}=\bigg{(}16\pi^{2}N\left(2G_{4}r_{h}-2X\left(\sqrt{2}\sqrt{f}s\sqrt{-X}\left(12F_{5}X-G_{5X}\right)+4F_{4}r_{h}X+2r_{h}G_{4X}\right)\right)\bigg{)}\delta r_{h}~{}.$ (52)
To derive this result, we have used the fact that $\delta f=-4\pi\delta r_{h}/\beta$, utilized the grand canonical ensemble (fixed $\beta$) and required $\delta\mathcal{I}_{E}=0$. Here $N$ will be a constant, which we may set equal to unity with a time re-parametrization and in the analytic examples that will follow we have set $N(r)=1$. All functions are evaluated at the event horizon of the black hole, and the entropy is obtained by integrating (52) with respect to the horizon and adjusting the integration constant that emerges so that the entropy vanishes in the absence of a black hole. Note here the absence of $G_{2},~{}G_{3}$ in this expression, as well as that parity breaking terms may contribute to the total entropy, as long as $X(r_{h})\neq 0$ and $\sqrt{f}\sqrt{-X}\left(12F_{5}X-G_{5X}\right)\neq 0$. When $F_{5}=G_{5X}/(12X)$ there is no contribution from the parity breaking terms and taking this into account, the no-ghost condition (9) implies that
@@
$\delta\mathcal{M}=\lim_{r\to\infty}\Bigg{(}\frac{8\pi N}{r}\left(2G_{4}r-2X\left(\sqrt{2}\sqrt{f}s\sqrt{-X}\left(12F_{5}X-G_{5X}\right)+4F_{4}rX+2rG_{4X}\right)\right)\Bigg{)}\delta m~{}.$ (55)
The mass may be obtained by evaluating the above limit and then integrating with respect to $m$. In the case where the parameter $m$ is not an independent integration constant, but has somehow been promoted to a constant of the theory appearing in the action, the mass and therefore the location of event horizon are both fixed by the theory. In such a non-physical scenario, the black holes will not radiate, thus having $T=0$ and according to the third law of thermodynamics will be described by zero entropy. Regarding the first law of thermodynamics, such theories will satisfy the following first law
@@
IV.2 Non-homogeneous Spacetimes
Here we will discuss the non-homogeneous spacetimes. In the non-homogeneous cases, one has to solve the theory, determine the behavior of $X$ and $f$ and then proceed with a calculation of the boundary terms. $X$ may contain the metric function $f$ and therefore the horizon radius (the black hole mass), and as a result, it is allowed to vary at the boundaries. Since we cannot have a general result as in the homogeneous case, we will discuss a very well-known black hole solution, that of [111, 114]. The thermodynamics of these solutions have been discussed in [111, 98, 115] following various methods. It has been observed that these methods disagree at some level. We will show that our general formalism leads to valid results for the mass and entropy of the black hole. This class of theories corresponds to Horndeski theory with $G_{2}(X)=X$ and $G_{4}(X)=m_{p}^{2}/2+zX/(2m_{p}^{2})$, closely following [111]. The black hole solution corresponds to a non-homogeneous spacetime described by
where $m$ is the ADM mass, while $X$ is dynamical and given by
$X(r)=\left(\frac{2z^{2}}{m_{p}^{6}r^{2}}+\frac{2z}{m_{p}^{4}}\right)^{-1}~{}.$ (59)
boundary of $\text{boundary of {\hbox{f}}}=\left(\frac{4\pi\beta r\sqrt{h(r)}m_{p}^{2}}{\sqrt{f(r)}}-\frac{4\pi\beta rz\sqrt{h(r)}X(r)}{\sqrt{f(r)}m_{p}^{2}}\right)\Bigg{|}_{r_{h}}^{\infty}.$ (60)
$\mathcal{M}=8\pi M~{},$ (62)
where $\zeta$ is an arbitrary (not determined from first principles) reference length scale. As we can see the entropy receives logarithmic corrections due to the non-minimal coupling between the scalar field and curvature. As we will see in the following examples the form of $G_{4}\sim 1+\alpha X$ yields logarithmic corrections in the framework of Horndeski theories. It is important to stress that the general result of (52) may hold in the case of non-homogeneous spacetimes when $\delta X=0$, however, one needs to compute the boundary term of $X$ and then decide whether it vanishes or not, since there is no way to realize a priori (before solving the field equations) whether the boundary of $X$ will survive, thus contributing to the entropy.
@@ V.1 General Relativity
In the case of GR we have $G_{4}=1$ while all other $G_{i},F_{i}$ are zero and $X=0$. Hence, for the Schwarzchild black hole $f=h=1-2m/r~{},N=1$ the boundary term (49) becomes
boundary of $\text{boundary of {\hbox{f}}}=4\pi\delta A~{}.$ (63)
In order to cancel this term (we require $\delta\mathcal{I}_{E}=0$), we have to introduce at the horizon via $\delta\mathcal{B}$ the boundary contribution
@@
boundary of $\text{boundary of {\hbox{f}}}=-16\pi\beta\delta m~{},$ (65)
@@ V.2 Shift and parity symmetric case
boundary of $\text{boundary of {\hbox{f}}}=8\pi r_{h}\beta N\delta f(G_{4}(X)-4F_{4}(X)X^{2}-2XG_{4X})~{}.$ (67)
@@
boundary of $\text{boundary of {\hbox{f}}}=8\pi f_{4}r_{h}\beta\delta f~{},$ (69)
@@
is saturated on the bound. As a result, any black hole solution arising in the context of shift and parity symmetric Horndeski theory that is described by $g_{tt}=1/g_{rr}$ will satisfy Bekenstein’s Area Law for the entropy [75] and preserve the NEC. This implies that there must be a deeper connection between the energy conditions and the entropy of a black hole [^1] This observation would not have been made possible, without the many discussions we have had with Panagiotis Dorlis and Sotirios-Neilos Vlachos.. Finally, the mass will be solely determined by the ADM mass, since the boundary term is identical to the one of the GR case. We also refer to Section III for a particular example. In the context of Horndeski cases, where the shift and parity-preserving term $F_{4}$ is vanishing, $F_{4}=0$, like [119, 39, 120], the choice of $G_{4}(X)=f_{4}+c_{1}\sqrt{X}$ will also preserve Bekenstein’s Area Law [75] for homogeneous spacetime metrics.
V.3 $4-$d Einstein-Scalar-Gauss-Bonnet-Gravity
@@
[…]
while the beyond Horndeski functions are both zero $F_{4}=0=F_{5}$ and the solution is a homogeneous black hole $h(r)=f(r)$ [27]. In particular, the solution reads ($s=-1$),
@@
boundary of $\text{boundary of {\hbox{f}}}=8\pi\beta\delta fN\left(r+4\sqrt{2}\alpha\sqrt{f}\sqrt{-X}-4\alpha rX\right)~{}.$ (74)
@@
where $\ell$ is an arbitrary (not determined from first principles) reference length scale and our result agrees with [21]. $\ell$ may be fixed so that extremal black holes will be described by zero entropy according to the third law (it has also been stated that extremal black holes have zero entropy [121, 122]. There is also the famous string theory result of non-zero entropy [123], however, in general, one should regard non-extreme and extreme black holes as qualitatively different objects). We note that only $G_{4}$ contributes to the entropy and leads to logarithmic corrections as we saw previously in the non-homogeneous spacetimes. Moving on to the contribution at infinity, we find the usual result of $\mathcal{B}(\infty)=16\pi m\beta$ which will give $\mathcal{M}=16\pi m$ for the conserved black hole mass.
@@ V.4 Parity-Breaking Theories Related to Lovelock Theory
This theory is particularly interesting due to its strong connections with higher-dimensional Lovelock gravity. In the special case $n=1$ with $\gamma c=-1$, where $G_{5}\sim\ln(X)$, the theory reduces to the one studied in [21], which emerges from the Kaluza-Klein reduction of the Gauss-Bonnet term in Lovelock gravity (see also [23] for an alternative derivation). Notably, this theory is equivalent to the previous example under the redefinition $\alpha\rightarrow-2\alpha$. More generally, when $\gamma c=-1$, the beyond Horndeski terms vanish, i.e., $F_{4}=0=F_{5}$, and the theory remains within the Horndeski class. In this scenario, the parameter $n+1$ can be directly associated with the order of the Lovelock term involved in the higher-dimensional reduction. For instance, the $n=2$ case, where $G_{4}=1+\alpha X^{2}$ and $G_{5}=X$, corresponds to the third-order Lovelock invariant $L_{3}$ which interestingly via Kaluza-Klein reduction to four dimensions gives the self-tuning Paul term [34] found in Fab 4 [124, 125]. Note that for $n=0$, the theory does not reduce to General Relativity, as $G_{4}=1$ but a kinetic term $G_{2}\sim X$ remains present. This arises due to the Kaluza-Klein reduction of the higher-dimensional Ricci scalar, which generically introduces additional scalar field dynamics in the four-dimensional effective theory. To recover General Relativity, one must take the limit $a\to 0$, which eliminates the scalar field contributions and restores the standard Einstein-Hilbert action.
@@
The solution for the aforementioned action functionals is a homogeneous black hole ($h(r)=f(r))$), where the kinetic term reads
@@
V.5 Three-dimensional Horndeski Theory

### v2

@@ IV General thermodynamical remarks
$\text{boundary of $X$}=$ $\bigg\{-4\pi\beta\sqrt{h}N\delta X\Big(\sqrt{2}s\sqrt{-X}\left(f\left(24X^{2}F_{5X}+60F_{5}X-2XG_{5XX}-3G_{5X}\right)+r^{2}G_{3X}+G_{5X}\right)$
$+4\sqrt{f}r\left(2X\left(2XF_{4X}+G_{4XX}\right)+G_{4X}\right)+32\sqrt{f}F_{4}rX\Big)\bigg\}\bigg|_{r_{h}}^{\infty}~.$ (48)
@@
$\text{boundary of $f$}=$ $\bigg(-4\pi\beta N\delta f\left(-2G_{4}r+2X\left(\sqrt{2}\sqrt{f}s\sqrt{-X}\left(12F_{5}X-G_{5X}\right)+4F_{4}rX+2rG_{4X}\right)\right)\frac{\sqrt{h}}{\sqrt{f}}\bigg)\bigg|_{r_{h}}^{\infty}~.$ (49)
@@
$\text{boundary of $X$}=\frac{4\pi\beta\delta X\sqrt{h}\left(fh^{\prime}-f^{\prime}h\right)\left(-2\sqrt{2}f^{3/2}s(-X)^{3/2}\left(12F_{5}X-G_{5X}\right)+G_{2}r^{3}+2G_{4}r\right)}{f^{3/2}X\left(rh^{\prime}+h\right)}~.$ (50)
Interestingly, for homogeneous metrics, which satisfy $h(r)=f(r)$, the first parenthesis vanishes, since $fh^{\prime}-f^{\prime}h=0$ so a boundary term for $X$ does not survive. In the non-homogeneous case, i.e. $h\neq f$, the boundary term of $X$ may vanish when the metric function $f$ satisfies
@@
The vanishing of the boundary term related to $X$ in the case of homogeneous spacetimes implies the absence of any conserved parameter related to the scalar field $\phi$, i.e a primary scalar hair, when $h(r)=f(r)$. However, in the case where the scalar field is linearly time-dependent, a nonzero conjugate momentum term arises when moving to the Hamiltonian formulation [120, 117], resulting to a non-zero boundary term for $X$. Hence, a linearly time-dependent scalar field is the only way that one can generate a black hole carrying primary scalar hair in the context of homogeneous metric, and the corresponding field has to have a non-zero conjugate momentum [31, 117], just like the case of electric fields in electromagnetism [125, 126, 87]. Regardless of the fact of the vanishing of the boundary term of $X$, a primary scalar hair has to be characterized by a non-trivial Noether current, which cannot be the case of purely radial scalar field, as it has been shown in [31].
In the following, we will set $N(r)=1.$, since the introduction of the lapse function $N$ was to be able to write the reduced Hamiltonian in a similar manner to the previous examples, and now there’s no reason to keep two different function for the same degree of freedom ($g_{tt}$).
@@ IV.1 Homogeneous Spacetimes
$\delta\mathcal{S}=\bigg(16\pi^{2}\left(2G_{4}r_{h}-2X\left(\sqrt{2}\sqrt{f}s\sqrt{-X}\left(12F_{5}X-G_{5X}\right)+4F_{4}r_{h}X+2r_{h}G_{4X}\right)\right)\bigg)\delta r_{h}~.$ (52)
To derive this result, we have used the fact that $\delta f=-4\pi\delta r_{h}/\beta$, utilized the grand canonical ensemble (fixed $\beta$) and required $\delta\mathcal{I}_{E}=0$. All functions are evaluated at the event horizon of the black hole, and the entropy is obtained by integrating (52) with respect to the horizon and adjusting the integration constant that emerges so that the entropy vanishes in the absence of a black hole. Note here the absence of $G_{2},~G_{3}$ in this expression, as well as that parity breaking terms may contribute to the total entropy, as long as $X(r_{h})\neq 0$ and $\sqrt{f}\sqrt{-X}\left(12F_{5}X-G_{5X}\right)\neq 0$. When $F_{5}=G_{5X}/(12X)$ there is no contribution from the parity breaking terms and taking this into account, the no-ghost condition (9) implies that
@@
$\delta\mathcal{M}=\lim_{r\to\infty}\Bigg(\frac{8\pi}{r}\left(2G_{4}r-2X\left(\sqrt{2}\sqrt{f}s\sqrt{-X}\left(12F_{5}X-G_{5X}\right)+4F_{4}rX+2rG_{4X}\right)\right)\Bigg)\delta m~.$ (55)
The mass may be obtained by evaluating the above limit and then integrating with respect to $m$. (In the case where the parameter $m$ is not an independent integration constant, but has somehow been promoted to a constant of the theory appearing in the action, the mass and therefore the location of event horizon are both fixed by the theory. In such a non-physical scenario, the black holes will not radiate, thus having $T=0$ and according to the third law of thermodynamics will be described by zero entropy.) Regarding the first law of thermodynamics, such theories will satisfy the following first law
@@ V.1 General Relativity
In the case of GR we have $G_{4}=1$ while all other $G_{i},F_{i}$ are zero and $X=0$. Hence, for the Schwarzchild black hole f=h=12m/r,N=1f=h=1-2m/r~,N=1 the boundary term (49) becomes
$\text{boundary of $f$}=8\pi\beta\delta fr\Big|_{r_{h}}^{\infty}~.$ (57)
In order to cancel this term (we require $\delta\mathcal{I}_{E}=0$), and according to our discussion so far we have to introduce at the horizon via $\delta\mathcal{B}$ the boundary contribution
@@
$\text{boundary of $f$}=-16\pi\beta\delta m~,$ (59)
@@ V.2 Shift and parity symmetric case
$\text{boundary of $f$}=8\pi r\beta\delta f(G_{4}(X)-4F_{4}(X)X^{2}-2XG_{4X})\Big|_{r_{h}}^{\infty}~.$ (61)
@@
$\text{boundary of $f$}=8\pi f_{4}r\beta\delta f\Big|_{r_{h}}^{\infty}~,$ (63)
@@
is saturated on the bound. As a result, any black hole solution arising in the context of shift and parity symmetric Horndeski theory that is described by $g_{tt}=1/g_{rr}$ will satisfy Bekenstein’s Area Law for the entropy [77] and preserve the NEC. This implies that there must be a deeper connection between the energy conditions and the entropy of a black hole [^2] This observation would not have been made possible, without the many discussions we have had with Panagiotis Dorlis and Sotirios-Neilos Vlachos.. It has also been proven [130] that the positivity of the energy density is required for the positivity of the entropy for non-extremal black holes in string-inspired models of gravity. Finally, the mass will be solely determined by the ADM mass, since the boundary term is identical to the one of the GR case. We also refer to Section III for a particular example. In the context of Horndeski cases, where the shift and parity-preserving term $F_{4}$ is vanishing, $F_{4}=0$, like [131, 40, 132, 133, 134], the choice of $G_{4}(X)=f_{4}+c_{1}\sqrt{X}$ will also preserve Bekenstein’s Area Law [77] for homogeneous spacetime metrics.
V.3 44-d Einstein-Scalar-Gauss-Bonnet-Gravity
@@
$G_{2}(X)=8\alpha X^{2}~,\hskip 5.69046ptG_{3}(X)=-8\alpha X~,\hskip 8.5359ptG_{4}(X)=1+4\alpha X~,\hskip 8.5359ptG_{5}(X)=-4\alpha\ln X~,$ (66)
while the beyond Horndeski functions are both zero $F_{4}=0=F_{5}$ and the solution is a homogeneous black hole $h(r)=f(r)$ [27]. In particular, the solution reads (s=1s=-1),
@@
$\text{boundary of $f$}=8\pi\beta\delta f\left(r+4\sqrt{2}\alpha\sqrt{f}\sqrt{-X}-4\alpha rX\right)\Big|_{r_{h}}^{\infty}~.$ (68)
@@
where $\ell$ is an arbitrary (not determined from first principles) reference length scale and our result agrees with [21]. $\ell$ may be fixed so that extremal black holes will be described by zero entropy according to the third law (it has also been stated that extremal black holes have zero entropy [135, 136]. There is also the famous string theory result of non-zero entropy [137], however, in general, one should regard non-extreme and extreme black holes as qualitatively different objects). We note that only $G_{4}$ contributes to the entropy and leads to logarithmic corrections as we will see in a following subsection, where we will discuss non-homogeneous spacetimes. Moving on to the contribution at infinity, we find the usual result of $\mathcal{B}(\infty)=16\pi m\beta$ which will give $\mathcal{M}=16\pi m$ for the conserved black hole mass. Finally, for interesting black-string solutions of this model, we refer to [138].
@@ V.4 Parity-Breaking Theories Related to Lovelock Theory
This theory is particularly interesting due to its strong connections with higher-dimensional Lovelock gravity. In the special case $n=1$ with $\gamma c=-1$, where $G_{5}\sim\ln(X)$, the theory reduces to the one studied in [21], which emerges from the Kaluza-Klein reduction of the Gauss-Bonnet term in Lovelock gravity (see also [23] for an alternative derivation). Notably, this theory is equivalent to the previous example under the redefinition $\alpha\rightarrow-2\alpha$. More generally, when $\gamma c=-1$, the beyond Horndeski terms vanish, i.e., $F_{4}=0=F_{5}$, and the theory remains within the Horndeski class. In this scenario, the parameter $n+1$ can be directly associated with the order of the Lovelock term involved in the higher-dimensional reduction. For instance, the $n=2$ case, where $G_{4}=1+\alpha X^{2}$ and $G_{5}=X$, corresponds to the third-order Lovelock invariant $L_{3}$ which interestingly via Kaluza-Klein reduction to four dimensions gives the self-tuning Paul term [35] found in Fab 4 [139, 140]. To recover General Relativity, one must take the limit $a\to 0$, which eliminates the scalar field contributions and restores the standard Einstein-Hilbert action.
@@
The solution for the aforementioned action functionals is a homogeneous black hole $h(r)=f(r)$, where the kinetic term reads
@@
V.5 Non-homogeneous Spacetimes
Here we will discuss the non-homogeneous spacetimes. In the non-homogeneous cases, one has to solve the theory, determine the behavior of $X$ and $f$ and then proceed with a calculation of the boundary terms. $X$ may contain the metric function $f$ and therefore the horizon radius (the black hole mass), and as a result, it is allowed to vary at the boundaries. Since we cannot have a general result as in the homogeneous case, we will discuss a very well-known black hole solution, that of [124, 142], which are solutions of the “kinetic coupling” model, where the Einstein tensor is coupled to the derivatives of the scalar field. For various interesting aspects of this theory we refer to [143, 144]. The thermodynamics of these solutions have been discussed in [124, 104, 145, 118] following various methods. It has been observed that these methods disagree at some level. This might be a computational issue, or something more profound might be at play. The Jacobson, Kang, Myers ambiguity [102] shows that such boundary-term ambiguities in the Wald formula do not affect the entropy of stationary black holes with a regular Killing horizon. However, in scalar–tensor theories these regularity conditions may fail, e.g. due to divergences of scalar derivatives, and the Wald formula may then become inapplicable. In fact, in [118] it was argued that in this case the temperature has to be altered in order for the first law of thermodynamics to hold, a scenario based on the fact that in modified gravity theories gravitons move on an effective metric. We will show that our general formalism leads to valid results for the mass and entropy of the black hole, while the temperature remains unaltered, as required by the Euclidean continuation and finite temperature systems. This class of theories corresponds to Horndeski theory with $G_{2}(X)=X$ and $G_{4}(X)=m_{p}^{2}/2+zX/(2m_{p}^{2})$, closely following [124]. The black hole solution corresponds to a non-homogeneous spacetime described by
where $m$ is the ADM mass, while $X$ is dynamical and given by
$X(r)=\left(\frac{2z^{2}}{m_{p}^{6}r^{2}}+\frac{2z}{m_{p}^{4}}\right)^{-1}~.$ (78)
$\text{boundary of $f$}=\left(\frac{4\pi\beta\delta fr\sqrt{h(r)}m_{p}^{2}}{\sqrt{f(r)}}-\frac{4\pi\beta\delta frz\sqrt{h(r)}X(r)}{\sqrt{f(r)}m_{p}^{2}}\right)\Bigg|_{r_{h}}^{\infty}.$ (79)
$\mathcal{M}=8\pi M~,$ (81)
where $\zeta$ is an arbitrary (not determined from first principles) reference length scale. As we can see the entropy receives logarithmic corrections due to the non-minimal coupling between the scalar field and curvature. As we saw previously in the subsection regarding the 4d4-d Einstein-Scalar-Gauss-Bonnet theory, the form of $G_{4}\sim 1+\alpha X$ yields logarithmic corrections in the framework of Horndeski theories. It is important to stress that the general result of (52) may hold in the case of non-homogeneous spacetimes when $\delta X=0$, however, one needs to compute the boundary term of $X$ and then decide whether it vanishes or not, since there is no way to realize a priori (before solving the field equations) whether the boundary of $X$ will survive, thus contributing to the entropy.
V.6 Three-dimensional Horndeski Theory